OpenStudy (anonymous):

The volume of a come is 120 cubic meters. What is the volume of a cylinder with the same height as the cone? A.30m B40m C360m D480m If you answer you get a medal

OpenStudy (anonymous):

its 360

OpenStudy (anonymous):

Vco = 1/3 x pi x r^2 x h Vcy = pi x r^2 x h Vco/Vcy = 1/3 120/Vcy = 1/3 Vcy = 360 cubic meter

OpenStudy (jdoe0001):

\(\bf \textit{volume of a cylinder}=\color{red}{\pi r^2 h}\\ \textit{volume of a cone}=\cfrac{\color{red}{\pi r^2 h}}{3}\)

OpenStudy (anonymous):

I'm the best response

OpenStudy (anonymous):

just sayin

OpenStudy (anonymous):

thanks @Nitotheburito @jdoe0001